# Rarefied neutral flow in orificed hollow cathodes

### v1

@@ Abstract
A parametric study is conducted to quantify the effect of the keeper electrode geometry on the neutral flow quantities (pressure, density, temperature) within orificed hollow cathodes, prior to cathode ignition. The keeper impinges directly on the flow out of the cathode orifice and its geometry strongly influences the product between the pressure in the orifice-keeper region and the distance between cathode and keeper, $P_{ko}\cdot D_{ko}$, a key parameter for successful cathode ignition. A representative cathode equipped with a keeper is simulated using the Direct Simulation Monte Carlo method. The numerical model is first validated with computational results from the literature, and a parametric study is then conducted. Parameters include the cathode pressure-diameter in the range of 1–5 Torr-cm and the following geometric ratios: cathode orifice-to-cathode inner radii in the range of 0.1–0.7, keeper orifice-to-cathode orifice radii in the range of 1–5, and keeper distance-to-cathode-orifice diameter, in the range of 0.5–10. It is found that, if both keeper and cathode have identical orifice radii, the flow remains subsonic in the cathode orifice-to-keeper region. In most cases, however, the flow becomes underexpanded and supersonic, and the static pressure within the cathode orifice-keeper region is, on average, 4% that of the upstream pressure value. The orifice-keeper region pressure increases with either (i) a decrease in the keeper orifice diameter or (ii) an increase in the distance between the cathode orifice plate and the keeper plate, in agreement with literature experimental data. Both trends are explained through control-volume-based conservation laws. The ratio of ignition-to-nominal mass flow rates is found to be in the range of 9–120, with a most probable value of 40, in agreement with experimental literature data. This suggests that heater-less cathode ignition at a minimum DC voltage may be achieved by increasing the input mass flow rate by a factor of 40.
The views expressed in this article are those of the authors and do not reflect the official policy or position of the Department of Defense, the U.S. Government, or any agency thereof.
The following article has been submitted to the Journal of Applied Physics. After it is published, it will be found at https://pubs.aip.org/aip/jap.
@@ I Introduction
A heater may be used to bring the emitter to electron emission temperatures prior to the start of the thermionic discharge. In this approach, the emitter is heated to high temperatures ($>1,200^{\circ}$), a neutral gas (e.g., xenon) is inserted in the cathode tube, and a discharge is established between the keeper and cathode. Alternatively, a DC, Paschen-like discharge may be established between the cathode and the keeper to start the cathode. Gas is inserted in the cathode tube, and a glow discharge produced by a high-voltage ($>100$ V) DC source effectively heats the cathode orifice and emitter. This approach is favored for low-to-medium amperage cathodes and circumvents the use of a heater: the heater and its associated power supply are oftentimes considered to be a single point-of-failure for space electric propulsion systems and require a dedicated power supply. The accurate placement and sizing of the keeper is critical for a reliable DC discharge: an optimal value of the pressure-distance product, $P_{ko}D_{ko}$, results in a minimum in the applied voltage required for the glow discharge, where $P_{ko}$ is the pressure of the gas in the orifice-keeper region, and $D_{ko}$ is the (axial) distance between the keeper and the cathode orifice plate. In the context of spacecraft electric propulsion, minimizing the applied voltage is desirable as it reduces the amount of on-board electric insulation required when using high-voltage power supplies.Lev et al. (2019) Although the keeper-orifice geometry is not identical to that of a Paschen discharge (e.g., the cathode orifice and keeper are not truly planar surfaces), the minimum pressure-distance product is typically $\sim$ 1–5 Torr-cm.Lev et al. (2019)
@@
In this work, we choose to revisit the study of Daykin-Iliopoulous et al. and explore, in depth, the effect of the keeper electrode on the fluid flow within orificed hollow cathodes that feature an enclosed keeper. Using a DSMC approach, we seek to identify which geometrical parameters most significantly impact the pressure-distance product in the orifice-keeper region of cathodes, provide broadly applicable design rules, and explain, physically, the observed trends. We first describe our numerical approach and validate our model based on reported literature data in Section II. We then perform a parameterized study based on set ranges of non-dimensional parameters that are reported in Section III. Scaled results are presented and discussed in Section IV.
@@ II.3.1 Physical domain and boundary conditions
We choose to apply diffuse boundary conditions throughout, with an accommodation coefficient of 0.8, and do not consider the influence of gas-surface interactions on our solution. Other studiesSharipov (2004) have shown, for example, that the flow rate through an orifice is not sensitive to the value of the accommodation coefficient for vessels separated by a thin orifice.
@@
The upstream surface features a subsonic inlet with a set pressure and temperature. A varying number of particles is injected to enforce the desired pressure and temperature. The total number of injected particles accounts for domain particles that exit the simulation domain through the upstream surface.
@@ II.3.2 Collision model
Particle-particle collisions are represented by the variable soft sphere (VSS) model. The corresponding VSS parameters are shown in Table 1 for xenon.
Table 1: VSS parameters for xenon (Ref. Bird, 1994, pp. 408,410)
Species $d_{\textrm{ref}}$(m) $\omega$ $T_{\textrm{ref}}$ $\alpha$
@@ II.3.4 Convergence and statistics
The domain is initialized with a non-zero number density that is computed based on a vacuum condition of 1 $\mu$Torr and a gas temperature of 300 K (i.e., $n_{n}=3.2\times 10^{16}$ m-3). Particles are drawn from a Maxwellian distribution with zero streaming velocity and the upstream temperature.
@@
Figure 3: Typical evolution of the total number of particles.
Relevant physical quantities are computed once steady-state is reached. Results are averaged over 10,000 time steps ($\sim$ 0.1 ms) to decrease the statistical noise inherent to particle simulations.
@@ II.5 Numerical validation
We validate the DSMC model with the test case described by Cao et al.,Cao et al. (2019) which features a computational domain without a keeper electrode. A fully coupled PIC/DSMC approach wherein both charged and neutral particles are simulated is beyond the scope of this article: we cannot capture plasma-neutral effects such as the increase in neutral particles temperature near the orifice due to the frequent charge-exchange collisions with the plasma ions.
The parameters describing the computational domain of the validation case are shown in Table 2. Using the perfect gas law and the provided upstream pressure and density (450 Pa and 1.3$\times$1022 m-3, respectively), the temperature of the injected gas is estimated to be 2,500 K. Unlike Cao et al., we do not rescale the heavy particle mass or domain size. For the set input upstream conditions, this results in an overestimation of the mass flow rate within the cathode. However, quantities such as the (static) pressure and the number density can be directly compared.
@@
$\dfrac{P_{\textrm{T,exit}}}{P_{\textrm{T,in}}}=\dfrac{1}{M_{in}}\left[\dfrac{2+\left(\gamma-1\right)M_{in}^{2}}{\gamma+1}\right]^{\dfrac{\gamma+1}{2\left(\gamma-1\right)}},$ (9)
@@
A Poiseuille flow approach is often used to estimate the pressure within the cathode.Goebel and Katz (2008); Mizrahi et al. (2012); Becatti, Conversano, and Goebel (2021); Becatti and Goebel (2022); Poli et al. (2023) We use the model from Goebel and KatzGoebel and Katz (2008) with the xenon viscosity from Stiel and ThodosStiel and Thodos (1961). The mass flow rate may be estimated by assuming a continuum flow and a choked orifice,
where $C_{d}$ is the discharge coefficient and $R_{g}$ the gas constant.
Figure 5 compares the pressure and the neutral density as obtained with our theoretical and numerical approaches to that of Cao et al. We also compare our results to the collisionless free-jet solution obtained by Cai and BoydCai and Boyd (2007) and to the results of the fluid solver ANSYS-CFX. The average orifice exit (static) temperature, gas velocity, and density scale required for the analytical model of Cai and BoydCai and Boyd (2007) are derived from our simulation. They are equal to 1,823 K, 447 m/s, and $5.0\times 10^{21}$ m-3, respectively. Reasonable agreement is obtained for all quantities for all numerical approaches. Because the flow is transitional in the orifice and near-orifice region (i.e., Knsimilar-$\sim$0.1–1), our numerical results differ from both the collisionless solution and the fluid solution. The DSMC and collisionless approaches yield the same density decay in the far field ($z\geq 12$ mm) where the Knudsen number increases. However, the fluid solver is unable to resolve this density decay. The Fanno flow model of the orifice yields the most accurate theoretical value. The Poiseuille flow approach overestimates the pressure and is, in general, inaccurate. As discussed elsewhere,Taunay, Wordingham, and Choueiri (2022a) this approach should not be used to estimate the pressure within cathodes as it is inapplicable to the flow regime in which cathodes operate: most of the assumptions required by this flow model are invalid in cathodes. For example, the flow is not fully developed in the orifice and compressibility effects are significant since the flow becomes sonic at the orifice exit.
Figure 5: On-axis (a) neutral pressure and (b) density for Cao et al.’s test case. All flow models start with the same value at $z=0$ mm. The isentropic and Fanno flow models are assumed to have the same static quantities at the orifice entrance. The error bars on the Poiseuille flow model indicate values of $C_{d}$ in the range of 0.1–1, with the symbol set at the pressure and density values for which $C_{d}=0.5$.
@@
Figure 6: Contours of (a) neutral density and (b) Knudsen number (normalized by the orifice diameter) for Cao et al.’s test case.
@@ III.1 Scope
Using Equation 10 and assuming $C_{d}=1$, the range of mass flow rates considered is approximately 0.28–13.5 mg/s (2.8–135 sccm of xenon). We limit our study to $d_{o}/d_{c}=$ 0.1 for $P_{T}d_{c}=5$ Torr-cm: for this particular pressure-diameter, the corresponding mass flow rate is above 34 mg/s (i.e., $\geq 300$ sccm) for $d_{o}/d_{c}=0.5$, which is much higher than typical cathode flow rates.
@@ IV.1 Flow structure in the presence of a keeper electrode
Accurately resolving the neutral flow field is critical for comprehensive plasma-neutral cathode simulations that use the DSMC method for the neutrals. The introduction of an enclosed keeper strongly affects the neutral flow downstream of the orifice. The flow structure differs from that of a simple expansion from the cathode orifice into vacuum (see, e.g., Figure 6(a)). In the presence of the keeper, the flow through the orifice first expands into a larger channel (i.e., the orifice-keeper region). After reaching the keeper, the flow undergoes a second expansion, into vacuum. Our numerical results reveal the existence of two main flow regimes: the flow may be either supersonic or subsonic in the orifice-keeper region.
@@ IV.1.1 Supersonic flow
In most cases, the pressure ratio $P_{ko}/P_{T}$ is such that the flow that leaves the cathode orifice is choked and supersonic, and the overall flow structure is similar to that reported by Titarev and ShakovTitarev and Shakhov (2014) for a composite pipe system composed of two pipes of increasing radii that exhaust into vacuum. Figures 11 through 14 illustrate the typical flow structure in the case where the orifice is choked. The flow expansion from the cathode orifice to the keeper-orifice region results in an underexpanded, rarefied, supersonic jet (Figures 111214). After exiting the orifice, the jet interacts with the keeper plate and a toroidal zone of recirculation appears in the region between the underexpanded jet and the top boundary (Figure 11). The Knudsen number in the keeper region is $\approx 0.1$ (Figure 13), indicating that the flow is in the transitional regime.
Refer to captionOrifice
@@
The exact structure of the underexpanded jet depends on the backing-to-upstream pressure ratio, and, therefore, the cathode geometry. The centerline Mach number, static pressure, and number density are shown in Figures 16 through 19 for $d_{o}/d_{c}=$ 0.1–0.7, respectively. We observe the following qualitative trends:

### v2

# Rarefied xenon flow in orificed hollow cathodes

@@ Abstract
A parametric study is conducted to quantify the effect of the keeper electrode geometry on the xenon neutral flow quantities within orificed hollow cathodes, prior to ignition. The keeper impinges directly on the flow out of the cathode orifice and its geometry influences the product between the pressure in the orifice-keeper region and the cathode-to-keeper distance. A representative cathode is simulated using the Direct Simulation Monte Carlo method. The numerical model is first validated with computational results from the literature. A parametric study is then conducted. Parameters include the cathode pressure-diameter in the range of 1–5 Torr-cm and the following geometric ratios (and ranges): cathode orifice-to-inner radii (0.1–0.7), keeper orifice-to-cathode orifice radii (1–5), and keeper distance-to-cathode-orifice diameter (0.5–10). It is found that, if both keeper and cathode have identical orifice radii, the flow remains subsonic in the orifice-to-keeper region. In most cases, however, the flow becomes underexpanded and supersonic, and the static pressure within the orifice-to-keeper region is, on average, 4% that of the upstream pressure value. The orifice-keeper region pressure increases with either a decrease in the keeper orifice diameter or an increase in the distance between cathode and keeper, in agreement with literature data. Both trends are explained through conservation laws. A statistical study of numerical results reveals that the ratio of ignition-to-nominal mass flow rates has a most probable value of 50, which suggests that heaterless cathode ignition at a minimum DC voltage may be achieved by increasing the input mass flow rate by a factor of 50.
The following article has been submitted to the Journal of Applied Physics.
@@ I Introduction
A heater may be used to bring the emitter to electron emission temperatures prior to the start of the thermionic discharge. In this approach, the emitter is heated to high temperatures ($>1,200^{\circ}$), a neutral gas (e.g., xenon) is inserted in the cathode tube, and a discharge is established between the keeper and cathode. Alternatively, a DC, Paschen-like discharge may be established between the cathode and the keeper to start the cathode. Gas is inserted in the cathode tube, and a glow discharge produced by a high-voltage ($>100$ V) DC source effectively heats the cathode orifice and emitter. This approach is favored for low-to-medium amperage cathodes and circumvents the use of a heater: the heater and its associated power supply are oftentimes considered to be a single point-of-failure for space electric propulsion systems. The accurate placement and sizing of the keeper is critical for a reliable DC discharge: an optimal value of the pressure-distance product, $P_{ko}D_{ko}$, results in a minimum in the applied voltage required for the glow discharge, where $P_{ko}$ is the pressure of the gas in the orifice-keeper region, and $D_{ko}$ is the (axial) distance between the keeper and the cathode orifice plate. In the context of spacecraft electric propulsion, minimizing the applied voltage is desirable as it reduces the amount of on-board electric insulation required when using high-voltage power supplies.Lev et al. (2019) Although the keeper-orifice geometry is not identical to that of a Paschen discharge (e.g., the cathode orifice and keeper are not truly planar surfaces), the minimum pressure-distance product is typically $\sim$ 1–5 Torr-cm.Lev et al. (2019)
@@
In this work, we choose to revisit the study of Daykin-Iliopoulous et al. and explore, in depth, the effect of the keeper electrode on the fluid flow within orificed hollow cathodes that feature an enclosed keeper. Using a DSMC approach, we seek to identify which geometrical parameters most significantly impact the pressure-distance product in the orifice-keeper region of cathodes, provide broadly applicable design rules, and explain, physically, the observed trends. We also compare DSMC results to that of a fluid solver in select cases to understand the limitations of continuum approaches applied to rarefied gas flows. We first describe our numerical approach and validate our model based on reported literature data in Section II. We then perform a parameterized study based on set ranges of non-dimensional parameters that are reported in Section III. Scaled results are presented and discussed in Section IV.
@@ II.3.1 Physical domain and boundary conditions
We choose to apply diffuse boundary conditions throughout with a choice of a constant accommodation coefficient for all solid surfaces. While the value of the accommodation coefficient may change appreciably depending on the surface material and roughness,Agrawal and Prabhu (2008); Selden et al. (2009) a sensitivity study of the solution on the accommodation coefficient is beyond the scope of this article: we do not consider the influence of gas-surface interactions on our solution and set the accommodation coefficient to a value of 0.8. While this value is 13% lower than the recommended value from Agrawal and Prabhu,Agrawal and Prabhu (2008) it is within the range of other reported data for xenon interaction with solid wallsSelden et al. (2009) and, therefore, represents a conservative value of the accommodation coefficient. Other studiesSharipov (2004) have shown that the flow rate through an orifice is not sensitive to the value of the accommodation coefficient for vessels separated by a thin orifice.
@@
The upstream surface features a subsonic inlet with a set pressure and temperature , as opposed to a set mass flow rate. A varying number of particles is injected to enforce the desired pressure and temperature. The total number of injected particles accounts for domain particles that exit the simulation domain through the upstream surface. The mass flow rate for the numerical experiments, $\dot{m}_{\textrm{exp}}$, may be retrieved by computing the flux of particles that cross a plane intersecting the cathode channel. The mass flow rate typically depends on the orifice geometry and may differ from the theoretical values that can be computed using a typical isentropic flow approach, $\dot{m}_{th}$. We define the discharge coefficient, $C_{d}$, as the ratio of the experimental-to-theoretical flow rates:
$C_{d}=\dfrac{\dot{m}_{\textrm{exp}}}{\dot{m}_{th}}.$ (1)
@@ II.3.2 Collision model
Particle-particle collisions are represented by the variable soft sphere (VSS) model, in which the relationship between the collision cross-section and the collision energy is represented by a power law. For this model, the collision deflection angle is $\chi=2\arccos((b/d)^{\left(1/\alpha\right)})$, where $b$ is the impact parameter. Consequently, the gas viscosity depends on the temperature through a power law: $\mu\propto T^{\omega}$. The model depends on four parameters:
  • $T_{\textrm{ref}}$: a reference temperature,
  • $d_{\textrm{ref}}$: the particle diameter for the collision cross-section at the reference temperature, $T_{\textrm{ref}}$,
  • $\omega$: the temperature dependence of the viscosity, and
  • $\alpha$: the scattering parameter.
The VSS parameters are shown in Table 1 for xenon.
Table 1: VSS parameters for xenon (Ref. Bird, 1994, pp. 408,410). $T_{\textrm{ref}}$, $d_{\textrm{ref}}$, $\omega$ and $\alpha$ are the reference temperature, the particle diameter for the collision cross section at the reference temperature ($T_{\textrm{ref}}$), the temperature dependence of the viscosity, and the scattering parameter, respectively.
Species $d_{\textrm{ref}}$ (m) $\omega$ $T_{\textrm{ref}}$ (K) $\alpha$
@@ II.3.3 Discretization
Figure 3: Typical evolution of the total number of particles.
@@ II.3.4 Convergence and statistics
The domain is initialized with a non-zero number density that is computed based on a vacuum condition of 1 $\mu$Torr and a gas temperature of 300 K (i.e., $n_{n}=3.2\times 10^{16}$ m-3). Particles are drawn from a Maxwellian distribution with zero streaming velocity and the upstream temperature. After initialization, additional particles are injected continuously at the inflow boundary described in Section II.C.1.c at a set temperature and pressure. Injected particles travel throughout the domain and interact with the different boundary conditions. They are removed from the domain if they reach the outflow (or the inflow) boundary.
@@
Relevant physical quantities (e.g., the gas temperature or the mass flow rate) are computed once steady-state is reached. Results are averaged over 10,000 time steps ($\sim$ 0.1 ms) to decrease the statistical noise inherent to particle simulations.
@@ II.4 Multi-level grid and load rebalance
Figure 5: Cathode computational representation for the numerical validation test case.
@@ II.5 Numerical validation
We validate the DSMC model with the test case described by Cao et al.,Cao et al. (2019) which features a computational domain without a keeper electrode. The computational domain for this test case is shown in Figure 5. A fully coupled PIC/DSMC approach wherein both charged and neutral particles are simulated is beyond the scope of this article: we cannot capture plasma-neutral effects such as the increase in neutral particles temperature near the orifice due to the frequent charge-exchange collisions with the plasma ions.
The parameters describing the computational domain of the validation case are shown in Table 2. Using the input parameters provided by Cao et al.Cao et al. (2019) for the upstream pressure and density (450 Pa and 1.3$\times$1022 m-3, respectively) and the perfect gas law, the temperature of the injected gas is found to be 2,500 K. Because we are considering a heavy, noble gas that is monatomic at high temperatures and low densities, the thermal de Broglie wavelength ($\mathcal{O}\left(10^{-12}\,\textrm{m}\right)$) is much smaller than the size of the domain ($\mathcal{O}\left(10^{-3}\,\textrm{m}\right)$): Boltzmann statistics still apply,McQuarrie (2000) and the perfect gas law is applicable. Unlike Cao et al., we do not rescale the heavy particle mass or domain size. For the set input upstream conditions, this results in an overestimation of the mass flow rate within the cathode. However, quantities such as the (static) pressure and the number density can be directly compared.
@@
$\dfrac{P_{\textrm{T,in}}}{P_{\textrm{T,exit}}}=\dfrac{1}{M_{in}}\left[\dfrac{2+\left(\gamma-1\right)M_{in}^{2}}{\gamma+1}\right]^{\dfrac{\gamma+1}{2\left(\gamma-1\right)}},$ (10)
@@
A Poiseuille flow approach is often used to estimate the pressure within the cathode.Goebel and Katz (2008); Mizrahi et al. (2012); Becatti, Conversano, and Goebel (2021); Becatti and Goebel (2022); Poli et al. (2023) We use the model from Goebel and KatzGoebel and Katz (2008) with the xenon viscosity from Stiel and ThodosStiel and Thodos (1961). The mass flow rate is computed from the simulation, as described in Section II.C.1.c, and is found to be 1.5 mg/s (15.3 sccm).
Figure 6 compares the pressure and the neutral density as obtained with our theoretical and numerical approaches to that of Cao et al. We also compare our results to the collisionless free-jet solution obtained by Cai and BoydCai and Boyd (2007) and to the results of the fluid solver ANSYS-CFX. The average orifice exit (static) temperature, gas velocity, and density scale required for the analytical model of Cai and BoydCai and Boyd (2007) are derived from our simulation. They are equal to 1,823 K, 447 m/s, and $5.0\times 10^{21}$ m-3, respectively. We estimate the average orifice gas temperature for the Poiseuille flow approach by assuming that the orifice is choked (i.e., the Mach number in the orifice is equal to one). The orifice static temperature used for the Poiseuille flow model is found to be equal to 1,873 K. Reasonable agreement is obtained for all quantities for all numerical approaches. Because the flow is transitional in the orifice and near-orifice region (i.e., Knsimilar-$\sim$0.1–1), our numerical results differ from both the collisionless solution and the fluid solution. The DSMC and collisionless approaches yield the same density decay in the far field ($z\geq 12$ mm) where the Knudsen number increases. However, the fluid solver is unable to resolve this density decay. The Fanno flow model of the orifice yields the most accurate theoretical value. The Poiseuille flow approach overestimates the pressure by 32% and 134% at the orifice entrance and exit, respectively. As discussed elsewhere,Taunay, Wordingham, and Choueiri (2022a) this approach should not be used to estimate the pressure within cathodes as it is inapplicable to the flow regime in which cathodes operate: most of the assumptions required by this flow model are invalid in cathodes. For example, the flow is not fully developed in the orifice and compressibility effects are significant since the flow becomes sonic at the orifice exit.
Figure 6: On-axis (a) neutral pressure and (b) density for Cao et al.’s test case. All flow models start with the same value at $z=0$ mm. The isentropic and Fanno flow models are assumed to have the same static quantities at the orifice entrance.
@@
Figure 7: Contours of (a) neutral density and (b) Knudsen number (normalized by the orifice diameter) for Cao et al.’s test case as obtained with our DSMC approach.
@@ III.1 Scope
The mass flow rate may be estimated by assuming a continuum flow and a choked orifice,
where $C_{d}$ is the discharge coefficient and $R_{g}$ the gas constant. Using Equation 11 and assuming $C_{d}=1$, the range of mass flow rates considered is approximately 0.28–13.5 mg/s (2.8–135 sccm of xenon). We limit our study to $d_{o}/d_{c}=$ 0.1 for $P_{T}d_{c}=5$ Torr-cm: for this particular pressure-diameter, the corresponding mass flow rate is above 34 mg/s (i.e., $\geq 300$ sccm) for $d_{o}/d_{c}=0.5$, which is much higher than typical cathode flow rates.
@@ IV.1 Flow structure in the presence of a keeper electrode
Accurately resolving the neutral flow field is critical for comprehensive plasma-neutral cathode simulations that use the DSMC method for the neutrals. The introduction of an enclosed keeper strongly affects the neutral flow downstream of the orifice. The flow structure differs from that of a simple expansion from the cathode orifice into vacuum (see, e.g., Figure 7(a)). In the presence of the keeper, the flow through the orifice first expands into a larger channel (i.e., the orifice-keeper region). After reaching the keeper, the flow undergoes a second expansion, into vacuum. Our numerical results reveal the existence of two main flow regimes: the flow may be either supersonic or subsonic in the orifice-keeper region, depending on the value of keeper-to-total pressure ratio, $P_{ko}/P_{T}$. The “keeper pressure”, $P_{ko}$, is defined as the average (static) pressure in the region bounded by the orifice plate and keeper entrance. Only the quiescent part of the flow (i.e., $v\approx 0$ m/s) is considered for the computation of $P_{ko}$. $P_{T}$ is the total upstream pressure. The pressure ratio we define is similar to the “ambient-to-total” (or “reservoir-to-total”) pressure ratio that is typically used to characterize nozzle flows.
@@ IV.1.1 Supersonic flow
In most cases, the pressure ratio $P_{ko}/P_{T}$ is such that the flow that leaves the cathode orifice is choked and supersonic, and the overall flow structure is similar to that reported by Titarev and ShakovTitarev and Shakhov (2014) for a composite pipe system composed of two pipes of increasing radii that exhaust into vacuum. Figures 12 through 15 illustrate the typical flow structure in the case where the orifice is choked. The flow expansion from the cathode orifice to the keeper-orifice region results in an underexpanded, rarefied, supersonic jet (Figures 121315). After exiting the orifice, the jet interacts with the keeper plate and a toroidal zone of recirculation appears in the region between the underexpanded jet and the top boundary (Figure 12). The Knudsen number in the keeper region is $\approx 0.1$ (Figure 14), indicating that the flow is in the transitional regime.
[…]
@@
The exact structure of the underexpanded jet depends on the backing-to-upstream pressure ratio, and, therefore, the cathode geometry. The centerline Mach number, static pressure, and number density are shown in Figures 18 through 21 for $d_{o}/d_{c}=$ 0.1–0.7, respectively. We observe the following qualitative trends:
@@
A statistical distribution of the discharge coefficient as obtained with the method outlined in Section II.C.1.c is shown in Figure 17 for the supersonic flow cases. The mass flow rate as obtained for the isentropic mass flow rate, $\dot{m}_{th}$, that is required to compute $C_{d}$ is given by
$\dot{m}_{th}=\dfrac{P_{T}\pi r_{o}^{2}}{\sqrt{T_{T}}}\sqrt{\dfrac{\gamma}{R_{g}}}\left(\dfrac{2}{\gamma+1}\right)^{\left(\gamma+1\right)/\left(\gamma-1\right)}.$ (14)
The average value and 95% confidence bounds for $C_{d}$ are 0.82, and 0.55 and 0.92, respectively. The study of the variation of $C_{d}$ with the geometric parameters (i.e., $d_{o}/d_{c}$, $d_{k}/d_{o}$, and $D_{ko}/d_{o}$) is beyond the scope of this article. For a cathode with a similar geometry as the one studied and for a set upstream pressure, a range of mass flow rate may be estimated with Equation 11 and $C_{d}$ in the 95% confidence interval. This approach may also be used to estimate the upstream (total) pressure of the gas, $P_{T}$, for a given mass flow rate.
Figure 17: Combined violin and box plot of the range of discharge coefficient, $C_{d}$. Each individual tick mark indicates the location of a data point.